OpenStudy (anonymous):

The volume of a cylinder is 24 ft3. The volume of a cube is 54 ft3. What is the difference between the volume of a cone that fits exactly inside the cylinder and the volume of a pyramid that fits exactly inside the cube? Provide an explanation for your answer to receive full credit.

OpenStudy (campbell_st):

the volume of the cone is 1/3 of the volume of the cylinder volume of the pyramid is 1/3 the volume of the cube then find the difference to get the answer

OpenStudy (anonymous):

So how would I solve that? @campbell_st

OpenStudy (campbell_st):

so simply divide the volume of the cylinder by 3... this will be the cone volume and then divide the volume of the cube by 3... this will be the pyramid volume then just subtract them

OpenStudy (anonymous):

Okay I understand so would this be right? 24/3 = 8 54/3 = 18 Then It'd be 24 - 8 = 16 and 54 - 18 = 46 @cambell_st

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so using the 1st part Cone V = 8ft^3 Pyramid V= 18 ft^3 the difference = 18 - 8 = 10 ft^3